OpenStudy (anonymous):

If f(x) = 2-x^2 + x^3 and f(a) = 2, then what are the the possible values of a? I dont get it, at all. Show workkk!

OpenStudy (anonymous):

2-a^2+a^3=2 -a^2+a^3=0

OpenStudy (anonymous):

so when a=0 , 0=0

OpenStudy (anonymous):

a^3=a^2 1=1

OpenStudy (anonymous):

so 0,1

OpenStudy (anonymous):

2-x^2 + x^3 = 2 -x^2 + x^3 = 0 x^2(-1 + x) = 0 x^2 = 0, x = 0 -1 + x = 0, x = 1 x = 0, 1

OpenStudy (anonymous):

a could also be 1

OpenStudy (anonymous):

Okay, thank you guys!(:

OpenStudy (anonymous):

yw kait